OpenStudy (anonymous):

Find the value for integral 1/2pi integral over limit 0 to infinity exp(-x2/8) dx

OpenStudy (jamesj):

First thing to know is the value of \[ \int_{-\infty}^{\infty} e^{-x^2} dx \] Do you know the value of this integral?

OpenStudy (jamesj):

My argument is that if you know the value of this integral, we can get to the value of your integral.

OpenStudy (anonymous):

I think we should use ILATE formula to integrate it

OpenStudy (jamesj):

I don't know what ILATE is but I'll tell you the value is \( \sqrt{\pi} \).

OpenStudy (jamesj):

That being the case, what then is the value of \[ \int_{-\infty}^{\infty} e^{-x^2/a^2} dx \] You can evaluate it by using a substitution u = x/a

OpenStudy (anonymous):

But in option a) 1, b) pi ,c) 2 and d)2pi are given and one of them is correct

OpenStudy (jamesj):

We haven't got to your integral yet. We're building up to it.

OpenStudy (anonymous):

Basically how use those integral symbol with limit i don't know. That's why I wrote the problem like that

OpenStudy (anonymous):

JamesJ, when your done with this question, can you please help me

OpenStudy (jamesj):

Your integral, \[ \frac{1}{2\pi} \int_0^{\infty} e^{-x^2/8} dx = \frac{1}{2\pi}\frac{1}{2} \int_{-\infty}^{\infty} e^{-x^2/8} dx \ \ \ \ \hbox{because the integrand is an even function} \]

OpenStudy (jamesj):

...an even function

OpenStudy (jamesj):

Therefore if you know how to find the value of \[ \int_{-\infty}^{\infty} e^{-x^2/a^2} dx \] you can find the value of your integral. So try the substitution I suggested.

OpenStudy (anonymous):

ok i'll try it out with this. thanks